A blue print of a house has a scale of 1.5 in : 5 ft. Would a ping pong table that is 9 feet by 5 feet fit in a space that has drawing dimensions of 3 inches and 4.8 inches? If so, how much space (in ft2) will remain after the table is put in the room?
Answer by Theo(13342)

the scale is 1.5 inches : 5 feet.

the ping pong table is 9 feet by 5 feet.

the space that the ping pong table will be placed in has drawing dimensions of 3 inches and 4.8 inches.

the ratio of real life to drawing dimensions is 5 feet / 1.5 inches.

the space in the drawing is 3 inches by 4.8 inches.

multiply by the scale factor and you get the space in real life is 3 inches * 5 feet / 1.5 inches by 4.8 inches * 5 feet / 1.5 inches.

that makes the space equal to 10 feet by 16 feet.

since the ping pong table is 5 feet by 9 feet, it will fit into the space very nicely.

the ping pong table takes 5 * 9 = 45 square feet.

the space takes 10 * 16 = 160 square feet.

after inserting the ping pong table into the space, there will be 160 - 45 = 115 square feet of space remaining
around it.